OpenStudy (anonymous):

PLEASE HELP ME IM SO SERIOUS PLEASEEE SOMEBODY For the data in the table, does y vary directly with x? If it does, write an equation for the direct variation. (Hint: you need to find the “k” value) x 20 30 40 y 2 10 18 PLEASE HELP MEEEEEE

OpenStudy (anonymous):

y=kx is the formula for direct variation. 2=20k divide by 20 k=1/10 10=30k divide by 30 k=1/3 18=40k divide by 40 k=9/20 So I would have to say no, it does not. Or else k would be constant between all of those.

OpenStudy (anonymous):

THANK YOUUUUU SO MUCHHHHHHHHHHHHHHH

OpenStudy (anonymous):

Not a problem. That's what we are here for. :3